# Towards a logic of affordances

### v1

@@ 8. Conclusion and future work
The above calls for addressing the following problem as well: are there “natural” axioms for affordances, for example, for knowledge spaces in learning theory
So far, on the level of operators, we have only used the definition of an affordance in terms of information systems as a source of intuition, allowing us to talk about equivalence classes of similar objects from three different sorts. Therefore, another task is to develop tools that allow for going down to the level of information systems and properties, and use this in reasoning with affordances.
@@ Acknowledgments
This research was partially funded by the National Science Center (Poland), grant number 2020/39/B/HS1/00216. Ivo Düntsch would like to thank Nicolaus Copernicus University in Toruń for supporting him through the “Excellence Initiative–Research University” program.

### v2

@@ 8. Conclusion and future work
The above calls for addressing the following problem as well: are there “natural” axioms for affordances, for example, for knowledge spaces in learning theory?
So far, on the level of operators, we have only used the definition of an affordance in terms of information systems as a source of intuition, allowing us to talk about equivalence classes of similar objects from three different sorts. Therefore, another task is to develop tools that allow going down to the level of information systems and properties, and use them in reasoning with affordances.
@@ Acknowledgments
This research was funded by the National Science Center (Poland), grant number 2020/39/B/HS1/00216. Ivo Düntsch would like to thank Nicolaus Copernicus University in Toruń for supporting him through the “Excellence Initiative–Research University” program.
The authors would like to thank the anonymous reviewers for their insightful and fair comments, which helped improve the paper.